OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. –2x – y = –7 2x – y = –7 2x – y = 11 –2x + y = –11

OpenStudy (anonymous):

Use the point slope formula: \[\large y - y_1 = m(x - x_1)\] \[y-5 = 2(x+1) \implies y - 5 = 2x + 2\] \[2x - y - -7\]

OpenStudy (anonymous):

Sorry, \[2x - y = -7\]

OpenStudy (anonymous):

thaankssss!!! (:

OpenStudy (anonymous):

Welcome dear..